# LEPTODERMIC CORRECTIONS TO THE TOV EQUATIONS AND NUCLEAR ASTROPHYSICS WITHIN THE EFFECTIVE SURFACE APPROXIMATION

### v1

@@ IV Discussions of the results
Figures 7 and 8 show the cuts of the contour plot of Fig. 6 at $A_{\rm S}\approx 1.15$ corresponding approximately to the typical NS mass $M=1.4M_{\odot}$ and $A_{\rm S}\approx 0.96$ for $M=2.0M_{\odot}$, respectively, as function of the radial coordinate $r$ (in km). These NS masses $M$ and approximately the radii by the order of the magnitude were found in the recent experiments presented in Refs. GR21 ; CC20 ; TR21 ; see also the comparison with the theoretical results in Ref. OL20 . We specify different values of the incompressibility parameter $\kappa$ from a weak (almost nuclear $\kappa=1$) to a strong value ($\kappa=10$) of the gravitational potential. As mentioned above, the value $\kappa=0$ is related to the pressure $p_{0}$ over zero order leptodermic parameter $a/R$; see Fig. 5. For example, one can consider stronger gravitational field up to the value $\kappa=10$ in Fig. 8. For nuclear physics, one has $\kappa\equiv K/(12b_{\rm V})=1.25$ for $K=240$ MeV, and $b_{\rm V}=16$ MeV (or $K=225$ MeV and $b_{V}=15$ MeV from Ref. AB14 ).
@@
Figure 9: The dimensionless pressures, $p$, Eq. (50), (thick solid black line for the surface component $p_{1}$ and thin red one for the total volume and surface quantity $p$) as functions of the dimensionless particle number density, $y$, (density $\rho$ in units of $\overline{\rho}$) for the solution (32) of leading order over the leptodermic parameter $a/R$, (a,c), and over the radial coordinate $r$, (b,d), for different dimensionless incompressibilites $\kappa=1$, (a,b), and $10$, (c,d), [see Eq. (49)], for the NS mass $1.4M_{\odot}$. Dashed (rare green $p_{1}$ and frequent blue $p$ ) lines show these pressures for the NS mass $2.0M_{\odot}$. The arrows show the ES value $y_{0}=1/3$ in (a,c) and the effective radius $R$ in (b,d).
@@
Figure 9 shows the EoS (a,c), $p(y)$, for small [$\kappa=1$, (a)] and large [$\kappa=10$, (c)] gravitational forces, and the same $p(y(r))$ (b,d) but with $y=y(r)$ for the vdW-Skyrme interaction, Eq. (32). Again, we compare two characteristic cases, small ($M=1.4M_{\odot}$, red solid) and relatively large ($M=2.0M_{\odot}$, blue frequent dashed) NS masses. The EoS pressures $p=p(y)$, Fig. 9(a,c), are increasing functions of the density $y$ while the pressures $p=p(y(r))$, (b,d), are decreasing as functions of the radial variable $r$ because of the leptodermic properties of the particle number density $y(r)$ for the vdW-Skyrme forces, Eq. (32). The surface ($S$) components, $p_{1}$, Eq. (48), are shown by black solid ($M=1.4M_{\odot}$) and green long dashed ($M=2.0M_{\odot}$) lines, respectively. The surface effects become much more significant with increasing gravitational parameter $\kappa$. These results for the pressure are basically in agreement with those presented in Fig. 4 for the macroscopic EoS. However, the surface correction $p_{1}$ leads to a more wide coordinate dependence of the total pressure $p$ near the ES (Fig. 9 (b,d)). As mentioned above, the NS masses $M$ and effective radii $R$ correspond largely to very accurate experimental data; see Refs. GR21 ; CC20 ; TR21 ; OL20 ). Notice also that the mass dependence is also essential for all results presented in Fig. 9 for the volume and surface contributions especially for a large gravitational field. We should emphasize that our results for the volume ($p_{V}$) contribution to the total pressure $p$ are in good agreement with the numerical calculations based on the TOV equations discussed in Ref. MS24 .
@@ V Conclusions
The effective surface approximation based on the leptodermic expansion over a small parameter - a ratio of the crust thickness $a$ to the effective radius $R$ of the system - is extended for the description of neutron star properties. The neutron star was considered as a dense finite liquid (including amorphous solid) system at the equilibrium. The mean gravitational potential $\Phi(\rho)$ was taken into account in the simplest form as expansion over powers of differences $\rho-\overline{\rho}$, where $\overline{\rho}$ is the saturation density, up to second order in terms of the particle separation energy and incompressibility. Taking into account the gradient terms of the energy density in a rather general form we analytically obtained the leading (over a small parameter $a/R\ll 1$) particle number density $\rho$. The density $\rho$ was derived as function of the normal-to-ES coordinate $\xi$ of the orthogonal local nonlinear-coordinate system $\xi,\eta$ through the effective surface. With its help, one finds the equation of state (EoS) for the pressure $P=P(\rho)$. The TOV equations were re-derived within the same ESA approximation taking into account the leptodermic surface corrections. Our results for the leading order volume pressure components are in good agreement with those of Tolman RT87 and numerical calculations MS24 . We obtained the first order leptodermic surface corrections owing to the Schwartzschild metric, and corresponding modified TOV equations, and have solved them analytically within the same order. These first order corrections, being of the order of $a/R$, are also proportional to the dimensionless incompressibility depending essentially on the gravitational interaction constant, along with the nuclear interaction part. The surface pressure corrections are significantly increased with increasing NS mass, and even more pronouncely, with the gravitational contribution to the total incompressibility.
As perspectives, one can generalize our analytical approach to the the many-component and rotating systems. It is especially interesting to extend this approach to take into account the symmetry energy of the isotopically asymmetric systems, and collective NS rotations. As this approach was formulated in the local nonlinear system of coordinates $\xi,\eta$ for deformed and superdeformed shapes of the NS surface, one can apply our method to the pair of neutron star pulsars at their large angular momenta.
@@ Appendix A Derivations for $\lambda_{1}$
Substituting Eqs. (24) and (26) into the second equation in Eq. (2), one has at zero order leptodermic parameter,
@@
Then, using this zero order equation, from Eq. (2), one obtains in the linear order in $\lambda_{1}$ ($G=c=1$),
@@
$\alpha(r)=re^{\lambda_{0}}\left(-8\pi\mathcal{E}_{0}+\frac{1}{r^{2}}\right)$
$=\frac{1}{r(1-\tilde{r}^{2})}-\frac{8\pi\mathcal{E}_{0}r}{1-\tilde{r}^{2}},$ (A4)
@@
Taking into account that $\alpha_{1}\propto\mathcal{E}_{1}\propto({\rm d}\rho/{\rm d}r)^{2}$, Eq. (A), at leading order of the leptodermic expansion, one can take the smooth coefficient in the integrand of Eq. (A11) off the integral at the ES, $r=R$. The reason is that the derivative square, $(\partial\rho/\partial r)^{2}$, is a bell-like function of $r$ with a relatively small width of the order of the crust thickness $a$. We assume here that $r$ is not close to the Schwarzschild radius $R_{\rm S}$, where the integrand in Eq. (A11) is singular. In line of Ref. MS24 , the remaining integration within the same precision can be taken analytically. Finally, for $C^{(\lambda)}_{1}(r)$ one finds ($r<R_{\rm S}$)
@@
where $\mathcal{D}_{1}$ is given by Eq. (29); see also Eq. (30) for the asymptote outside and Eq. (31) [in terms of $\mathcal{Y}(y)$, Eq. (III.2)] inside of the system. The effective radius $R$ is presented in units of $R_{\rm S}$ and $c=G=1$, for shortness. From Eqs. (A12) and (A10), one arrives at the final expression (28) for $\lambda_{1}$.

### v2

@@ IV Discussions of the results
Figures 7 and 8 show the cuts of the contour plot of Fig. 6 at $A_{\rm S}\approx 1.15$ corresponding approximately to the typical NS mass $M=1.4M_{\odot}$ and $A_{\rm S}\approx 0.96$ for $M=2.0M_{\odot}$, respectively, as function of the radial coordinate $r$ (in km). These NS masses $M$ and approximately the radii by the order of the magnitude were found in the recent experiments presented in Refs. GR21 ; CC20 ; TR21 ; see also the comparison with the theoretical results in Ref. OL20 . We specify different values of the incompressibility parameter $\kappa$ from a weak (almost nuclear $\kappa=1$) to a strong value ($\kappa=10$) of the gravitational field. As mentioned above, the value $\kappa=0$ is related to the pressure $p_{0}$ for the zero order leptodermic parameter $a/R$; see Fig. 5. For example, one can consider stronger gravitational field up to the value $\kappa=10$ in Figs. 7 and 8. For nuclear physics, one has $\kappa\equiv-K/(12b_{\rm V})=1.25$ for $K=240$ MeV, and $b_{\rm V}=-16$ MeV (or $K=225$ MeV and $b_{V}=-15$ MeV from Ref. AB14 ).
@@
Figure 9: The dimensionless pressures, $p$, Eq. (50), (thick solid black line for the surface component $p_{1}$ and thin red one for the total volume and surface pressure $p$) as functions of the dimensionless particle number density, $y$, (density $\rho$ in units of $\overline{\rho}$) for the solution (32) of leading order over the leptodermic parameter $a/R$, (a,c), and over the radial coordinate $r$, (b,d), for different dimensionless incompressibilites $\kappa=1$, (a,b), and $10$, (c,d), [see Eq. (49)], for the NS mass $1.4M_{\odot}$. Dashed (rare green $p_{1}$ and frequent blue $p$ ) lines show these pressures for the NS mass $2.0M_{\odot}$. The arrows show the ES value $y_{0}=1/3$ in (a,c) and the effective radius $R$ in (b,d).
@@
Figure 9 shows the EoS (a,c), $p(y)$, for small $\kappa=1$, (a)] and large $\kappa=10$, (c)] gravitational fields, and the same $p(y(r))$ (b,d) but with $y=y(r)$ for the vdW-Skyrme interaction, Eq. (32). Again, we compare two characteristic cases, small ($M=1.4M_{\odot}$, red solid) and relatively large ($M=2.0M_{\odot}$, blue frequent dashed) NS masses. The EoS pressures $p=p(y)$, Fig. 9(a,c), are increasing functions of the density $y$ while the pressures $p=p(y(r))$, (b,d), are decreasing as functions of the radial variable $r$ because of the leptodermic properties of the particle number density $y(r)$ for the vdW-Skyrme forces, Eq. (32). The surface ($S$) components, $p_{1}$, Eq. (48), are shown by black solid ($M=1.4M_{\odot}$) and green long dashed ($M=2.0M_{\odot}$) lines, respectively. The surface effects become much more significant with increasing incompressibility parameter $\kappa$, that corresponds to a stronger gravitation field. These results for the pressure are basically in agreement with those presented in Fig. 4 for the macroscopic EoS. However, the surface correction $p_{1}$ leads to a more wide coordinate dependence of the total pressure $p$ near the ES (Fig. 9 (b,d)). As mentioned above, the NS masses $M$ and effective radii $R$ correspond largely to very accurate experimental data; see Refs. GR21 ; CC20 ; TR21 ; OL20 ). Notice also that the mass dependence is also essential for all results presented in Fig. 9 for the volume and surface contributions especially for a large gravitational field. We should emphasize that our results for the volume ($p_{V}$) contribution to the total pressure $p$ are in good agreement with the numerical calculations based on the TOV equations discussed in Ref. MM24 .
@@ V Conclusions
The effective surface approximation based on the leptodermic expansion over a small parameter - a ratio of the crust thickness $a$ to the effective radius $R$ of the system - is extended for the description of neutron star properties. The neutron star was considered as a finite dense and perfect liquid (including amorphous solid) system at the equilibrium. The mean gravitational field $\Phi(\rho)$ was taken into account in the simplest form as expansion over powers of differences $\rho-\overline{\rho}$, where $\overline{\rho}$ is the saturation density, up to second order in terms of the particle separation energy and incompressibility. Taking into account the gradient terms of the energy density in a rather general form we analytically obtained the leading (over a small parameter $a/R\ll 1$) particle number density $\rho$. The density $\rho$ was derived as function of the normal-to-ES coordinate $\xi$ of the orthogonal local nonlinear-coordinate system $\xi,\eta$ through the effective surface. With its help, one finds the macroscopic equation of state (EoS) for the pressure $P=P(\rho)$. The TOV equations were re-derived within the same ESA approximation taking into account the leptodermic surface corrections. Our results for the leading order volume pressure components are in good agreement with those of Tolman RT87 and numerical MM24 calculations. We obtained the first order leptodermic surface corrections owing to the Schwartzschild metric, and corresponding modified TOV equations, and have solved them analytically within the same order. These first order corrections, being of the order of $a/R$, are also proportional to the dimensionless incompressibility depending essentially on the gravitational interaction constant, along with the nuclear interaction part. The surface pressure corrections are significantly increased with increasing NS mass, and even more pronouncely, with the gravitational contribution to the total incompressibility.
As perspectives, one can generalize our analytical approach to the the many-component and rotating systems. It is especially interesting to extend this approach to take into account the symmetry energy of the isotopically asymmetric systems, and the collective NS rotations within the Kerr metric approach at small rotational frequencies and deformations. As this approach was formulated in the local nonlinear system of coordinates $\xi,\eta$ for deformed shapes of the NS surface, one can apply our method to the pair of rotating neutron star pulsars.
@@ Appendix A Derivations for $\lambda_{1}$
Substituting Eqs. (24) and (26) into the second equation in Eq. (II), one has at zero order leptodermic parameter ($G=c=1$),
@@
Then, using this zero order equation, from Eq. (II), one obtains in the linear order over $\lambda_{1}$,
@@
$\alpha(r)=re^{\lambda_{0}}\left(-8\pi\mathcal{E}_{0}+\frac{1}{r^{2}}\right)$ (A4)
$=\frac{1}{r(1-\tilde{r}^{2})}-\frac{8\pi\mathcal{E}_{0}r}{1-\tilde{r}^{2}},$
@@
Taking into account that $\alpha_{1}\propto\mathcal{E}_{1}\propto({\rm d}\rho/{\rm d}r)^{2}$, Eq. (A), at leading order of the leptodermic expansion, one can take the smooth coefficient in the integrand of Eq. (A11) off the integral at the ES, $r=R$. The reason is that the derivative square, $(\partial\rho/\partial r)^{2}$, is a bell-like function of $r$ with a relatively small width of the order of the crust thickness $a$. We assume here that $r$ is not close to the Schwarzschild radius $R_{\rm S}$, where the integrand in Eq. (A11) is singular. In line of Ref. MM24 , the remaining integration within the same precision can be taken analytically. Finally, for $C^{(\lambda)}_{1}(r)$ one finds ($r<R_{\rm S}$)
@@
where $\mathcal{D}_{1}$ is given by Eq. (29); see also Eq. (30) for the asymptote outside and Eq. (31) [in terms of $\mathcal{Y}(y)$, Eq. (III.2)] inside of the system. The effective radius $R$ is presented in units of $R_{\rm S}$. From Eqs. (A12) and (A10), one arrives at the final expression (28) for $\lambda_{1}$.